OpenStudy (anonymous):

convert each equation to standard form and determine the locations of the center vertices co vertices and foci 2x^2+3y^2-8x+6y+5=0

OpenStudy (agreene):

if by standard form you mean: y=ax^b+cx^d...+x+n then that is going to be really annoying and messy from the looks of things O.o

OpenStudy (agreene):

Click show steps (they solve by completing the square, which is what I was thinking about, lol) http://www2.wolframalpha.com/input/?i=2x%5E2%2B3y%5E2-8x%2B6y%2B5%3D0+solve+for+y

OpenStudy (anonymous):

standard form as in x^2/a^2+ Y^2/b^2 =1 (thats for horizontal ellipse

OpenStudy (agreene):

\[\frac{1}{2}(x-2)^2+\frac{1}{2}(y+1)^2=1\] Is the closest to that type of thing I can see.